# Predicting Multi-Agent Specialization via Task Parallelizability

### v1

@@ Abstract
Multi-agent systems often rely on specialized agents with distinct roles rather than general-purpose agents that perform the entire task independently. However, the conditions that govern the optimal degree of specialization remain poorly understood. In this work, we propose that specialist teams outperform generalist ones when environmental constraints limit task parallelizability—the potential to execute task components concurrently. Drawing inspiration from distributed systems, we introduce a heuristic to predict the relative efficiency of generalist versus specialist teams by estimating the speed-up achieved when two agents perform a task in parallel rather than focus on complementary subtasks. We validate this heuristic through three multi-agent reinforcement learning (MARL) experiments in Overcooked-AI, demonstrating that key factors limiting task parallelizability influence specialization. We also observe that as the state space expands, agents tend to converge on specialist strategies, even when generalist ones are theoretically more efficient, highlighting potential biases in MARL training algorithms. Our findings provide a principled framework for interpreting specialization given the task and environment, and introduce a novel benchmark for evaluating whether MARL finds optimal strategies.
multi-agent reinforcement learning, collaboration, efficiency, specialization, policy diversity
@@ 1 Introduction
Collaboration enables multi-agent systems to achieve goals that exceed the capabilities of a single agent. Advances in artificial intelligence (AI) have enhanced the capabilities of multi-agent systems to complete complex, coordinated tasks, including automating warehouses (Canese et al., 2021), sustainable resource management (Perolat et al., 2017), and simulated social interactions (Park et al., 2023). However, efficient task allocation remains a key challenge.
Existing work across multiple fields, including multi-agent reinforcement learning (MARL), cognitive science, and robotics, often assumes that specialization is universally desirable (Padgham & Winikoff, 2002; Zhu & Zhou, 2008; Charakorn et al., 2020; Wang et al., 2020a, b; Li et al., 2021; Bettini et al., 2023; Juang et al., 2024; Bettini et al., 2024; Swanson et al., 2024). For example, fine dining restaurants rely on specialists—the saucier prepares sauces, the rôtisseur roasts, and so on—who each perform distinct yet complementary subtasks. Intuitively, specialization narrows action spaces (Wang et al., 2020b), reduces ambiguity about what agents should do next (Goldstone et al., 2024), and eases individual cognitive loads (Griffiths, 2020). This frames task allocation as an optimization problem, focusing on identifying roles and assigning agents accordingly (Wang et al., 2020a).
However, human teams demonstrate that specialization is not always optimal. Fine dining chefs coordinate to produce complex meals, but at the cost of high training overhead, rigidity, and failure risk if a key specialist is unavailable (Ben-Oren et al., 2023; Vélez et al., 2024). In contrast, generalist fry cooks perform the same subtasks in parallel, such as grilling and assembling, maximizing adaptability and throughput. Examples of specialist and generalist strategies are depicted in Figure 1. Furthermore, although AI agents also face coordination challenges (Shoham & Leyton-Brown, 2008; Wu et al., 2021), it remains unclear whether the constraints that make specialization advantageous for people also apply. With agents becoming increasingly capable we must revisit this assumption and ask: under what conditions, if any, is specialization truly optimal?
@@
Figure 1: Specialists vs. generalists. (A) Specialist teams tend to focus on non-overlapping subtasks. In Overcooked, an example of this is Agent 1 putting onions and tomatoes in the pot, while Agent 2 fetches a bowl and serves the soup. In highly specialized teams, agents’ high-level action distributions are completely distinct, resulting in a Jensen-Shannon divergence ($\text{JSD}(P_{1}||P_{2})$) of 1.0. Example specialist video. (B) In generalist teams, both agents perform all subtasks. In Overcooked, this may look like Agents 1 and 2 independently making their own soups in parallel. In generalist teams, agents’ action distributions are the same, resulting in $\text{JSD}(P_{1}||P_{2})=0.0$. Example generalist video.
We propose that a task’s parallelizability—the potential for its independent subtasks to be executed simultaneously—limits the benefits of generalist teams. Drawing from the distributed computing literature (Almasi & Gottlieb, 1994; McCool et al., 2012; Vélez et al., 2023), we use Amdahl’s Law to determine when multi-agent systems should favor generalists or specialists. Amdahl’s Law predicts that there are diminishing returns to increasing the number of processors performing a task in parallel, based on the portion of the task that must be performed serially (Amdahl, 1967; Hill & Marty, 2008; Cassidy & Andreou, 2011). Applied to multi-agent systems, this places an upper bound on performance gains achievable from generalist agents working in parallel. When spatial and resource constraints limit parallel execution, agents should improve efficiency by specializing in complementary subtasks. Using a simple, interpretable model, we provide a novel framework for evaluating the trade-offs between specialist and generalist strategies.
To test this prediction, we develop a theoretical framework for comparing generalist and specialist efficiency and apply it to three multi-agent experiments in Overcooked-AI. Experiment 1 shows that key model variables influence specialization, with a moderate negative correlation between predicted task parallelizability and specialization, though larger state spaces complicate this relationship. Experiment 2 controls for these confounds, demonstrating our model’s high accuracy in predicting specialization. Experiment 3 finds that larger state spaces promote specialization by increasing coordination challenges. Our findings suggest that specialization is driven by task parallelizability and environment, enabling the design of layouts that encourage or inhibit policy diversity. More broadly, our framework highlights the value of using insights from distributed computing to improve the interpretability of multi-agent systems.
2 Theoretical Framework
We introduce a general framework for evaluating task parallelizability and predicting emergent specialization in multi-agent systems.
@@ 2.1 $N$-Player Markov Game
Following Carroll et al. (2019) and Ruhdorfer et al. (2024), we formalize a cooperative decentralized Markov decision process (MDP), defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{N},A,\mathcal{T},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ is a finite set of states, and $\mathcal{N}$ is a set of $N$ agents. $A$ is the joint action space, where each agent $i\in\mathcal{N}$ selects an action $a_{i}\in A$ to form a joint action $a=(a_{1},a_{2},...,a_{N})$. $\mathcal{T}:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition function which defines the probability distribution over next states given the current state and joint action. Finally, $\mathcal{R}:\mathcal{S}\times A\to\mathbb{R}$ is the shared reward function. The objective of each agent is to maximize its expected discounted return $\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\big{]}$, where $\gamma\in[0,1)$ is a discount factor and $r(s_{t},a_{t})$ is the shared reward received by all agents at time $t$. While agents share the same reward, each optimizes its own cumulative return independently.

### v2

@@ Abstract
When should we encourage specialization in multi-agent systems versus train generalists that perform the entire task independently? We propose that specialization largely depends on task parallelizability: the potential for multiple agents to execute task components concurrently. Drawing inspiration from Amdahl’s Law in distributed systems, we present a closed-form bound that predicts when specialization improves performance, depending only on task concurrency and team size. We validate our model on two standard MARL benchmarks that represent opposite regimes— StarCraft Multi-Agent Challenge (SMAC, unlimited concurrency) and Multi-Particle Environment (MPE, unit-capacity bottlenecks)—and observe close alignment between the bound at each extreme and an empirical measure of specialization. Three follow-up experiments in Overcooked-AI demonstrate that the model works in environments with more complex spatial and resource bottlenecks that allow for a range of strategies. Beyond prediction, the bound also serves as a diagnostic tool, highlighting biases in MARL training algorithms that cause sub-optimal convergence to specialist strategies with larger state spaces.
@@ 1 Introduction
Collaboration enables multi-agent systems to achieve goals that exceed the capabilities of a single agent. Advances in artificial intelligence (AI) have enhanced the capabilities of multi-agent systems to complete complex, coordinated tasks, including automating warehouses [1], sustainable resource management [2], and simulated social interactions [3]. The prevailing assumption in multi-agent reinforcement learning (MARL), cognitive science, and robotics is that specialization is universally desirable [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. For example, fine dining restaurants rely on specialists—the saucier prepares sauces, the rôtisseur roasts, and so on—who each perform distinct yet complementary subtasks. Intuitively, specialization narrows action spaces, reduces ambiguity about what agents should do next [14], and eases individual cognitive loads [15]. Recent algorithms such as ROMA and RODE rely on this assumption to treat task allocation as an optimization problem, focusing on learning roles and constraining agents to specific subtasks [8, 7]. However, in both human and MARL teams, specialization introduces trade-offs. Fine dining chefs produce complex meals via rigid role divisions, but the system is brittle: training costs are high, and service halts if one specialist fails or takes longer than the others [16, 17]. In contrast, generalist fry cooks, who all grill, assemble, and plate, maximize throughput by performing subtasks in parallel. While MARL agents do not fatigue or quit like humans, they do face coordination challenges, policy divergence, and retraining costs when roles are overly fixed [18, 19]. This raises a key question: what task and environmental constraints make specialization truly optimal, and when are generalists better?
@@
Figure 1: Specialists vs. generalists. (A) Specialist teams focus on non-overlapping subtasks, resulting in high Jensen-Shannon divergence (SI) between agents’ action distributions. Example video. (B) In generalist teams, both agents perform the task independently, resulting in low SI. Video.
We propose that the degree to which a task can be executed in parallel (or its parallelizability) directly governs the effectiveness of generalist teams. When spatial or resource bottlenecks force agents to wait, the benefits of independently performing a task in parallel vanish and specialist policies become more efficient. Drawing from distributed computing [20, 21], we adapt Amdahl’s Law to treat agents as processors and subtasks as units of work that may or may not be parallelizable [22, 23, 24]. This yields a simple analytic bound on the speed-up achievable by generalist teams as a function of task structure. When the parallelizability of the overall task falls below team size, specialization is advantageous.
We show that a task’s parallelizability provides a simple prediction for when specialist policies will emerge in multi-agent systems given a team size and environment. In Experiment 1, we apply the model to two edge-case MARL benchmarks: SMAC, where generalists dominate due to unlimited concurrency, and MPE, where unit-capacity bottlenecks force agents to specialize. In Experiments 2 and 3, we extend to more complex spatial and resource dynamics in Overcooked-AI, demonstrating a more fine-grained relationship between task parallelizability and specialization. Experiment 4 shows that larger state spaces promote deviations from model predictions, highlighting biases in existing MARL training regimes. Together, these findings suggest that specialization is not obligatory for successful teams, and is affected by bottlenecks and coordination structure. Our framework enables the design of environments that encourage or inhibit policy diversity and illustrates how tools from distributed systems can illuminate multi-agent settings.
2 Task Parallelizability Framework
In many multi-agent settings, the marginal benefit of adding agents is limited not only by training but also structural constraints on the task and environment. Quantifying task parallelizability answers two key questions: (1) When should we encourage behavioral diversity and specialization during training, and (2) When parallelism is limited, which dimension— team size $N$, subtask allocation $f_{i}$, spatial capacity $C^{s}$, or resource capacity $C^{r}$— is responsible for preventing speed-up? The key result (§2.4) shows that when overall parallelizability $S(N,B)$ drops below team size, specialization becomes strictly more efficient. Sections 3–5 demonstrate how to measure concurrency constraints for various MARL applications and validate the prediction empirically.
@@ 2.1 $N$-Player Markov Game
Following Albrecht et al. [25], we formalize a cooperative decentralized Markov decision process (Dec-MDP), defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{N},A,\mathcal{T},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ is a finite set of states, and $\mathcal{N}$ is a set of $N$ agents. $A$ is the joint action space, where each agent $i\in\mathcal{N}$ selects an action $a_{i}\in A$ to form a joint action $a=(a_{1},a_{2},...,a_{N})$. $\mathcal{T}:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition function which defines the probability distribution over next states given the current state and joint action. Finally, $\mathcal{R}:\mathcal{S}\times A\to\mathbb{R}$ is the shared reward function. Each agent’s objective is to maximize its expected discounted return $\mathbb{E}\big[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\big]$, where $\gamma\in[0,1)$ is a discount factor and $r(s_{t},a_{t})$ is the shared reward received by all agents at time $t$. Agents share the same reward but optimize their returns independently.